A flywheel initially at rest accelerated uniformly at 6 rads-2. Calculate the angular velocity after 20 s and its angular displacement.


1
Expert's answer
2021-04-12T06:56:09-0400

The angular velocity for the constanly accelerating (from the rest) body is given as follows:


"\omega = \alpha t"

where "\alpha = 6rad/s^2" is the angular acceleration, and "t = 20s" is the time. Thus, obtain:


"\omega = 6\cdot 20 = 120\space rad/s"

The angular displacement is:


"\varphi = \dfrac{\alpha t^2}{2} = \dfrac{6\cdot 20^2}{2} = 1200 \space rad"

Answer. Angular velocity: 120 rad/s, angular displacement: 1200 rad.